# Sup-norm bounds for Jacobi cusp forms

### v2

@@ 4. Sup-norm bounds for Jacobi cusp forms
For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be a Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, which is normalized with respect to the Petersson inner product. Then, we have the bound
(61) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|^{2}_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{2}}\,m^{2+\epsilon}\big{)},$
where the implied constant depends only on the choice of $\epsilon>0$.
@@
Substituting the decomposition (13) of the Jacobi form $f(\tau,z)$ into its pointwise Petersson norm and applying the Cauchy–Schwartz inequality, we compute
$\|f(\tau,z)\|^{2}_{\mathrm{Pet}}$ $=\bigg{|}\sum_{\mu=0}^{2m-1}\varphi_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z)\bigg{|}^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}$
$\leq\bigg{(}\sum_{\mu=0}^{2m-1}|\varphi_{\mu}(\tau)|^{2}\,\eta^{k-1/2}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}|\vartheta_{\mu,m}(\tau,z)|\,\eta^{1/2}\,e^{-\frac{4\pi my^{2}}{\eta}}\bigg{)}$
$=\bigg{(}\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\bigg{)}.$
Thus
(62) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|^{2}_{\mathrm{Pet}}\leq\sup_{\tau\in\mathbb{H}}\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\cdot\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}.$
Since $f$ is normalized with respect to the Petersson inner product, we have from (15) that
(63) $\langle f,f\rangle_{\mathrm{Pet}}=\frac{1}{\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]}\sum_{\mu=0}^{2m-1}\langle\varphi_{\mu},\varphi_{\mu}\rangle_{\mathrm{Pet}}=1.$
Combining (63) with the bound (35) given in Theorem 3.5, we find the bound
(64) $\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}=\sum_{\mu=0}^{2m-1}\langle\varphi_{\mu},\varphi_{\mu}\rangle_{\mathrm{Pet}}\,\frac{\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}}{\langle\varphi_{\mu},\varphi_{\mu}\rangle_{\mathrm{Pet}}}\leq C\,\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]\,k^{\frac{3}{2}},$
where $C$ is a positive constant depending on $\Gamma_{0}$. Recalling now that
(65) $[\Gamma_{0}:\Gamma_{1}]=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]=O_{\epsilon}\big{(}m^{1+\epsilon}\big{)},$
where the implied constant depends only on the choice of $\epsilon>0$, we derive from the bounds (64) and (65), the bound
(66) $\sup_{\tau\in\mathbb{H}}\sum_{\mu=0}^{2m-1}\|\varphi_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{2}}\,m^{\frac{3}{2}+\epsilon}\big{)},$
where the implied constant depends only on the choice of $\epsilon>0$. The claimed bound (61) finally follows from (62) by combining the bound (66) with the bound (59) established in Proposition 4.3. ∎
@@
The bound (61) is polynomial in $k$ and $m$, and thus improves W. Kohnen’s bound (1), which is exponential in $k$. Comparing our bound with the one obtained by P. Anamby and S. Das in [AD23], there is an improvement with regard to the polynomial growth in $k$, while the polynomial growth in $m$ is slightly worse.
@@ Acknowledgements
Finally, the authors express their gratitude to Edgar Assing, whose observations have helped to improve our initial bound (3) by a factor of $\sqrt{m}$.

### v3

@@ 4. Sup-norm bounds for Jacobi cusp forms
For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be an $L^{2}$-normalized Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$. Then, we have the $L^{\infty}$-norm bound
(49) $\|\phi\|^{2}_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}=O_{\Gamma_{0},\epsilon}\big{(}k^{2}\,m^{\frac{7}{2}+\epsilon}\big{)},$
where the implied constant depends on $\Gamma_{0}$ and the choice of $\epsilon>0$.
@@
Substituting the decomposition (13) of the Jacobi form $\phi(\tau,z)$ into its pointwise Petersson norm and applying the Cauchy–Schwartz inequality, we find the estimate
$\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}$ $=\bigg{|}\sum_{\mu=0}^{2m-1}h_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z)\bigg{|}^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}$
$\leq\bigg{(}\sum_{\mu=0}^{2m-1}|h_{\mu}(\tau)|^{2}\,\eta^{k-\frac{1}{2}}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}|\vartheta_{\mu,m}(\tau,z)|^{2}\,\eta^{\frac{1}{2}}\,e^{-\frac{4\pi my^{2}}{\eta}}\bigg{)}$
$=\bigg{(}\sum_{\mu=0}^{2m-1}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\bigg{)}\bigg{(}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}^{2}\bigg{)}.$
Thus, we arrive by means of Lemma 4.1 at the bound
$\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}$ $\leq\sum_{\mu=0}^{2m-1}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\cdot\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}^{2}$
$\leq 2m\sum_{\mu=0}^{2m-1}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\,\eta^{\frac{1}{2}}+O\bigg{(}m^{\frac{1}{2}}\,\sum_{\mu=0}^{2m-1}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\bigg{)},$
from which we derive, after recalling the comment at the beginning of this section,
$\|\phi\|^{2}_{L^{\infty}}$ $=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}=\sup_{\begin{subarray}{c}\tau\in\mathcal{F}_{\Gamma_{0}}\\ z\in E_{\tau}\end{subarray}}\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}$
(50) $\leq 2m\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\big{(}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\,\eta^{\frac{1}{2}}\big{)}+O\bigg{(}m^{\frac{1}{2}}\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\bigg{)}.$
In order to be able to apply Lemma 4.2 and Theorem 3.5 to the two summands in (50), respectively, we need to $L^{2}$-normalize the modular forms $h_{\mu}$ ($\mu=0,\ldots,2m-1$) under consideration. To do so, we observe that formula (15) in conjunction with the $L^{2}$-normalization of $\phi$ gives
(51) $\|\phi\|^{2}_{L^{2}}=\frac{1}{\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]}\sum_{\mu=0}^{2m-1}\|h_{\mu}\|^{2}_{L^{2}}=1.$
An application of Lemma 4.3 to (51) then gives
$\sum_{\mu=0}^{2m-1}\|h_{\mu}\|^{2}_{L^{2}}=\sqrt{4m}\,[\Gamma_{0}:\Gamma_{1}]=O_{\epsilon}\big{(}m^{\frac{5}{2}+\epsilon}\big{)}$
with an implied constant depending on the choice of $\epsilon>0$. With regard to the first summand in (50), we thus obtain by Lemma 4.2 the bound
$2m\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\big{(}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}\,\eta^{\frac{1}{2}}\big{)}$ $=2m\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\bigg{(}\frac{\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}}{\|h_{\mu}\|^{2}_{L^{2}}}\,\eta^{\frac{1}{2}}\bigg{)}\|h_{\mu}\|^{2}_{L^{2}}$
(52) $=O_{\Gamma_{0}}\big{(}2mk^{2}\big{)}\sum_{\mu=0}^{2m-1}\|h_{\mu}\|^{2}_{L^{2}}=O_{\Gamma_{0},\epsilon}\big{(}k^{2}\,m^{\frac{7}{2}+\epsilon}\big{)}$
with an implied constant depending on $\Gamma_{0}$ and the choice of $\epsilon>0$. With regard to the second summand in (50), we derive from Theorem 3.5 the bound
$m^{\frac{1}{2}}\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}$ $=m^{\frac{1}{2}}\sum_{\mu=0}^{2m-1}\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\bigg{(}\frac{\|h_{\mu}(\tau)\|^{2}_{\mathrm{Pet}}}{\|h_{\mu}\|^{2}_{L^{2}}}\bigg{)}\|h_{\mu}\|^{2}_{L^{2}}$
(53) $=O_{\Gamma_{0}}\big{(}m^{\frac{1}{2}}k^{\frac{3}{2}}\big{)}\sum_{\mu=0}^{2m-1}\|h_{\mu}\|^{2}_{L^{2}}=O_{\Gamma_{0},\epsilon}\big{(}k^{\frac{3}{2}}\,m^{3+\epsilon}\big{)}$
with an implied constant depending on $\Gamma_{0}$ and the choice of $\epsilon>0$. The bounds (52) and (53) complete the proof of the theorem. ∎
@@
Comparing our bound with the one obtained by P. Anamby and S. Das in [AD23], we obtain the same polynomial growth in $k$, while the polynomial growth of their bound is better by a factor of $m^{\frac{3}{4}}$.